# How In-Context Learning Emerges from Training on Unstructured Data: On the Role of Co-Occurrence, Positional Information, and Noise Structures

### v1

@@ Abstract
Large language models (LLMs) like transformers have impressive in-context learning (ICL) capabilities; they can generate predictions for new queries based on input-output sequences in prompts without parameter updates. While many theories have attempted to explain ICL, they often focus on structured training data similar to ICL tasks, such as regression. In practice, however, these models are trained in an unsupervised manner on unstructured text data, which bears little resemblance to ICL tasks. To this end, we investigate how ICL emerges from unsupervised training on unstructured data. The key observation is that ICL can arise simply by modeling co-occurrence information using classical language models like continuous bag of words (CBOW), which we theoretically prove and empirically validate. Furthermore, we establish the necessity of positional information and noise structure to generalize ICL to unseen data. Finally, we present instances where ICL fails and provide theoretical explanations; they suggest that the ICL ability of LLMs to identify certain tasks can be sensitive to the structure of the training data. [^1] Software that replicates the empirical studies can be found at https://github.com/yixinw-lab/icl-unstructured.
@@ Introduction
Large language models (LLMs) such as transformers demonstrate impressive in-context learning (ICL) abilities [Brown et al., 2020]: without updating parameters, they can identify tasks and generate predictions from prompts containing input-output examples. For example, given the prompt dog anjing, cat kucing, lion singa, elephant, a well-trained LLM should detect the English-to-Indonesian pattern and predict gajah, the Indonesian translation for elephant, as the most likely next token. The emergence of ICL is surprising for at least two reasons. First, LLMs are trained from unstructured natural language data in an unsupervised manner through next-token prediction. Second, the training data of LLMs likely does not include sentences that resemble typical ICL prompts, i.e., of the form $c_{1}d_{1}\cdots c_{K}d_{K}$, where $(c_{k},d_{k})$ represents a known input-output pair.
Many efforts have sought to understand ICL from various theoretical and empirical perspectives; see related work in Section 5. Some studies (e.g., Akyürek et al. [2022], Von Oswald et al. [2023], Dai et al. [2023], Zhang et al. [2024], Ahn et al. [2024]) expanded Garg et al.’s [2022] regression formulation and attributed transformers’ ICL ability to gradient descent. Other studies (e.g., Wang et al. [2023], Zhang et al. [2023], Chiang and Yogatama [2024]) adopted a Bayesian perspective, building upon Xie et al.’s [2021] argument that ICL performs implicit Bayesian inference. While these connections are theoretically intriguing, they do not fully capture the actual ICL phenomenon: ICL arises from training on unstructured natural language data that are distinct from ICL prompts.
Figure 1: This paper aims to understand how in-context learning (ICL) emerges from pretraining on unstructured natural language data. In Section 2, we show that ICL can arise merely through modeling co-occurrence information using continuous bag of words (CBOW). Violet represents relationship-specific noise tokens. In Section 3, we establish the necessity of positional information and blocked noise structures for certain ICL tasks. Violet represents noise tokens. In Section 4, we present two scenarios where ICL unexpectedly fails and provide theoretical explanations, highlighting the importance of training data structure in enabling the ICL ability of language models.
This work. We study how ICL emerges from pretraining on unstructured natural language data. Throughout the paper, we focus on two types of ICL tasks. The first type involves known input-output pairings that frequently occur together in a sentence, such as (country)-(capital) and (English word)-(Indonesian translation). The second type involves recognizable patterns that may not commonly co-occur in a sentence, such as (word)-(first letter).
For the first task (see left of Figure 1), we examine cases where the training sentences contain one or two distinct input-output relationship types. We also consider more realistic scenarios where some input-output pairs do not always occur together and two types of relationships can co-occur in a single sentence. We prove that, in most cases, ICL is achievable by only modeling co-occurrence information using continuous bag of words (CBOW) [Mikolov et al., 2013], a language model from the pre-transformer era. Further, we conduct prompting and synthetic data experiments that support the conclusion that ICL may indeed arise from co-occurrence information.
@@
For the second task (see middle of Figure 1), we investigate cases where the training sentences contain one or two distinct patterns, as well as a more realistic scenario where noise tokens are present. We prove that positional information and blocked noise structures (e.g., pqrs in Figure 1) are crucial for the success of ICL. This observation aligns with Chen et al.’s [2024b] empirical finding that parallel structures in pretraining data support ICL. Moreover, we find that learned positional embeddings generally perform better, except in noisy scenarios where the noises are not clustered in blocks.
@@
Finally, we present two scenarios where ICL unexpectedly fails regardless of model architectures; see right of Figure 1. In the first scenario (left example), both the training data and test prompts follow repeating patterns across blocks, but the pattern being repeated in the test data differs from that in the training data. In the second scenario (right example), training sentences contain known input-output pairs but only at fixed locations. These findings, along with their empirical and theoretical explanations, underscore that LLMs may require specific structures in the pretraining data to exhibit ICL ability.
@@
Summary of contributions. In this paper, we (1) theoretically and empirically show that ICL can arise from merely modeling co-occurrence patterns using CBOW, (2) prove that, in other instances, ICL requires modeling positional information and blocked noise structures, and (3) present scenarios where ICL fails, highlighting the crucial role of training data structure for the emergence of ICL.
@@
In this section, we focus on in-context learning (ICL) tasks involving pairings that commonly co-occur within training sentences. To motivate the discussion, we revisit the (English word)-(Indonesian translation) example in Section 1. Below we perform a simple experiment with ChatGPT 3.5 [OpenAI, 2022]. The model is given prompts of the following form:
Provide the most plausible next token to complete this sentence (only the answer). Even if the sentence does not make sense, please complete it as best as you can: dog anjing, cat kucing, lion singa, [word]
We take turns replacing [word] with elephant, tiger, soon, and main. For the first two options, ChatGPT 3.5 correctly outputs gajah and harimau, their respective Indonesian translations. However, it does not provide the correct outputs for the latter two: it follows soon with lebih baik beri makanan haiwan! (better feed the animals!) and main with bola (ball). [^2] In Indonesian, main means play, main bola means play soccer, and mainan means toy. A similar pattern is observed with LLaMA 2 [Touvron et al., 2023], which produces the correct translations the first two words but incorrectly continues the last two words with to-be-published and an11footnotemark: 1, respectively.
If ICL stems from the ability of LLMs to recognize consistent mappings in test prompts, these models should be equally likely to produce the correct answer for any given [word], irrespective of its relevance to the in-context examples. However, this experiment demonstrates that this is not the case; in Section 2.4, we also present two experiments involving countries, US states, and their capital cities. This naturally raises the question: Does/can ICL arise from modeling co-occurrence information using a simple model like continuous bag of words (CBOW) [Mikolov et al., 2013]?
ICL via CBOW. We prove that, for certain tasks, ICL is achievable by modeling co-occurrence information between pairs of tokens (regardless of their positions) using CBOW. We utilize a variant of CBOW where each center word is modeled conditional on all other words in a sentence. We associate each word $w$ with their center and context embeddings $u_{w}$ and $v_{w}$ of the same dimension. Given a sentence $x_{1}x_{2}\cdots x_{I}$, the $i$-th word is distributed conditional on the other words in the sentence as follows:
$p(x_{i}=k\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{-i})\propto\exp\left(\frac{u_{k}^{\top}\sum_{j\neq i}v_{x_{j}}}{I-1}\right).$
The $u_{w}$’s and $v_{w}$’s are learned by minimizing the sum of the cross-entropy losses across all sentences and positions.
@@
Roadmap of Section 2. In Section 2.1, we begin by considering a simple ICL task of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, where $(c_{i},d_{i})$ represents a known pairing (e.g., a country and its capital city) and $i_{1},i_{2},\cdots,i_{\ell+1}$ are all distinct. The focus is to investigate whether a trained CBOW model can correctly output $d_{i_{\ell}}$. We also explore two other scenarios: ICL tasks of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$ in Section 2.2 (two connected relationships), as well as $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell-1}}d_{i_{\ell-1}}c_{i_{\ell}}$ and $e_{i_{1}}f_{i_{1}}\cdots e_{i_{\ell}}f_{i_{\ell}}e_{i_{\ell+1}}$ (two disconnected relationships) in Section 2.3. Sections 2.4 and 2.5 conclude with prompting and synthetic data experiments that provide support to the theory.
2.1 In-context learning on single-relationship tasks
@@
We investigate ICL in single-relationship tasks that take the form of $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, where $(c_{i},d_{i})$’s denote known pairings such as countries and their capital cities; there is only one type of relationship between $c_{i}$ and $d_{i}$. The vocabulary consists of $c_{1:K},d_{1:K},r_{1:L}$, where $r_{i}^{\prime}s$ represent other words (e.g., stop words). We first introduce Theorem 1, which states that ICL can arise if each sentence consists of exactly one $(c_{i},d_{i})$ pair, as long as the number of in-context examples ($\ell$) is not too large. To simplify calculations, we replace the cross-entropy loss with the squared loss. This involves removing the softmax activation and comparing the outputs against the one-hot encoding of the target words. The proof of Theorem 1 is in Appendix A.
Theorem 1 (ICL on single-relationship tasks).
@@
Let $K,L\geq S\geq 3$. Suppose we have infinitely many training sentences of length $S$, each comprising exactly one $(c_{i},d_{i})$ pair and $S-2$ distinct $r_{i}$’s. We train a CBOW model with the squared loss and a sufficiently large embedding dimension on these sentences. Given a prompt $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ with distinct $i_{k}$’s, the model predicts $d_{i_{\ell+1}}$ if and only if
@@
As an example, when each training sentence contains exactly one country-capital pair (i.e., $(c_{i},d_{i})$), Theorem 1 says a trained CBOW model will correctly predict $d_{i_{\ell+1}}$ (i.e., the capital city of $c_{i_{\ell+1}}$) given an ICL prompt of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, provided that the prompt length ($12\ell+$ 1) is not too large. Intuitively, this behavior is due to the presence of $c_{i_{\ell+1}}$ in the ICL prompt, which leads the model to correctly predict $d_{i_{\ell+1}}$ due to the frequent occurrences of the pair $(c_{i_{\ell+1}},d_{i_{\ell+1}})$ in the training data. If we let $L\rightarrow\infty$ and fix $K$ and $S$, the condition in Theorem 1 becomes $2\ell+1<K(S-1)^{2}/(S-2)^{2}$. This inequality trivially holds if the prompt length is set to be $S-1$ to match the training sentences.
Moreover, the same analysis applies when each sentence consists of exactly two (instead of one) different $(c_{i},d_{i})$ pairs. Again, letting $L\rightarrow\infty$ and fixing $K$ and $S$, the model correctly predicts $d_{i_{\ell+1}}$ given the same ICL prompt if and only if $2\ell+1<\frac{K(K-2)(S-1)^{2}}{(K-2)(S-2)(S-4)-K}$. This upper bound is strictly larger than $K(S-1)^{2}/(S-2)^{2}$: when each sentence contains exactly two $(c_{i},d_{i})$ pairs, ICL under the squared loss occurs for longer prompts.
Experiments. To empirically verify Theorem 1 and its generalizations, we conduct experiments using the cross-entropy loss with $S=8$, $K=10$, $L=20$, and $\ell=3$. We explore multiple $(p_{0},p_{1},p_{2})$ values, where $p_{k}$ denotes the probability of having exactly $k$ pairs of $(c_{i},d_{i})$ in the sentence. For each $(p_{0},p_{1},p_{2})$ triple, we introduce a more realistic setting where $c_{i}$ and $d_{i}$ do not always appear together by considering its corrupted version. In this setup, each $(c_{i},d_{i})$ pair has a 25% chance of being replaced with $(c_{i},r_{j})$ and a 25% chance of being replaced with $(d_{i},r_{j})$, for some $j\in[L]$.
Table 1: ICL on various single-relationship tasks, averaged over 10 repetitions, demonstrates stable, good performance across embedding dimensions ($d_{E}$), as Theorem 1 suggests. The corrupted setting also demonstrates excellent ICL ability under certain scenarios.
@@
Table 1 displays the average accuracy for each scenario, calculated over 10 repetitions. Notably, when $(p_{0},p_{1},p_{2})$ is $(0,1,0)$ or $(0,0,1)$, ICL under the cross-entropy loss achieves zero accuracy, in contrast to perfect accuracy with the squared loss as shown in Theorem 1. We believe this difference in accuracy is an artifact of the loss functions used, although its relevance is limited by the fact that, in reality, it is unlikely for every sentence to contain at least one $(c_{i},d_{i})$ pair. On the other hand, perfect ICL performance is observed in other settings (e.g., when the training sentences contain either zero, one, or two $(c_{i},d_{i})$ pairs) in both the clean and corrupted scenarios.
2.2 In-context learning on dual-connected-relationship tasks
In Section 2.1, we discussed the case where the training sentences contain one relationship, namely $(c_{i},d_{i})$’s. We now explore ICL on dual-connected-relationship tasks, where the two types of relationships are connected and denoted by $(c_{i},d_{i})$ and $(c_{i},e_{i})$: $c_{i}$ might represent a country, $d_{i}$ its capital city, and $e_{i}$ its currency. The vocabulary comprises $c_{1:K},d_{1:K},e_{1:K},r_{1:L}$, where $r_{i}$’s represent other words. The corresponding ICL tasks thus take the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$, where the model is expected to output $d_{i_{\ell+1}}$ and $e_{i_{\ell+1}}$. This involves task selection as the model should use the in-context examples to infer the task. We first present Theorem 2, which states that a trained CBOW model can perform task selection if each sentence contains exactly two distinct $(c_{i},d_{i})$ pairs or two distinct $(c_{i},e_{i})$ pairs with uniform probability. Its proof is in Appendix B.

### v2

# From Unstructured Data to In-Context Learning: Exploring What Tasks Can Be Learned and When

@@ Abstract
Large language models (LLMs) like transformers demonstrate impressive in-context learning (ICL) capabilities, allowing them to make predictions for new tasks based on prompt exemplars without parameter updates. While existing ICL theories often assume structured training data resembling ICL tasks (e.g., x-y pairs for linear regression), LLMs are typically trained unsupervised on unstructured text, such as web content, which lacks clear parallels to tasks like word analogy. To address this gap, we examine what enables ICL in models trained on unstructured data, focusing on critical sequence model requirements and training data structure. We find that many ICL capabilities can emerge simply from co-occurrence of semantically related word pairs in unstructured data; word analogy completion, for example, can provably arise purely through co-occurrence modeling, using classical language models like continuous bag of words (CBOW), without needing positional information or attention mechanisms. However, positional information becomes crucial for logic reasoning tasks requiring generalization to unseen tokens. Finally, we identify two cases where ICL fails: one in logic reasoning tasks that require generalizing to new, unseen patterns, and another in analogy completion where relevant word pairs appear only in fixed training positions. These findings suggest that LLMs’ ICL abilities depend heavily on the structural elements within their training data. [^1] Software that replicates the empirical studies can be found at https://github.com/yixinw-lab/icl-unstructured. Details on implementation, experiments and data sets are provided in Appendix J.
@@ Introduction
Large language models (LLMs) such as transformers demonstrate remarkable in-context learning (ICL) abilities [10]: without any parameter updates, they can recognize tasks and generate predictions from prompt examples. For instance, given the prompt dog anjing, cat kucing, lion singa, elephant, a well-trained LLM should detect the English-to-Indonesian pattern in the prompt and predict gajah—the Indonesian translation for elephant—as the most likely next token. The ICL capabilities of LLMs are surprising for two main reasons. First, these models are trained in an unsupervised manner on unstructured natural language data through next-token prediction, without any loss function specifically designed for ICL. Second, the training data for LLMs likely lacks sequences resembling typical ICL prompts, i.e., of the form $c_{1}d_{1}\cdots c_{K}d_{K}$, where $(c_{k},d_{k})$’s represent word pairs with specific semantic relationships.
Many efforts have sought to understand ICL from theoretical and empirical perspectives, e.g., gradient descent in regression and Bayesian inference. While insightful, these analyses often rely on structured training data that mirrors ICL tasks. For instance, they train on sequences of x-y pairs from various linear regression tasks and test on similar data. In practice, however, LLMs are trained in an unsupervised manner on unstructured text data, such as web content, which bears little resemblance to typical ICL tasks like word analogy. Consequently, these analyses may only partially capture the complexities of ICL.
This work. We investigate common ICL tasks to identify what tasks can be learned in context by a model trained on unstructured data. Specifically, we examine essential components of sequence modeling that enable in-context learning, along with requirements on the unstructured training data.
The first set of (theoretical and empirical) results focuses on ICL for word analogy completion using frequently co-occurring tokens [10, 50]. This task involves identifying relationships between word pairs, such as (country)-(capital) and (English word)-(Indonesian translation), then applying the same relationship to complete a sequence. For this task (see left of Figure 1), we explore cases where training sentences contain one or two types of word pairs with distinct semantic relationships. We prove that, in most cases, ICL can arise by simply modeling word co-occurrence using classical (pre-transformer) language models like continuous bag of words (CBOW) [36], without needing positional information or attention mechanisms.
@@
The second set of results involves ICL for logic reasoning tasks that require recognizing patterns that do not commonly co-occur in a sentence, such as (word)-(first letter) [62, 13]. For this task (see middle of Figure 1), we investigate scenarios where training sentences contain one or two distinct patterns, as well as a more realistic scenario where nuisance tokens are present. We prove that positional information and blocked nuisance structure (e.g., pqrs in Figure 1) are crucial for the success of ICL in these tasks. This finding aligns with Chen et al.’s [2024b] observation that parallel structures in pre-training data support ICL. We also find that learned positional embeddings generally perform better, except in scenarios where the nuisance tokens are not clustered in blocks.
Figure 1: This paper identifies essential components for in-context learning (ICL) from pre-training on unstructured natural language data. Left sub-panels, right sub-panels, and boxed letters denote NLP examples, our abstractions, and expected outputs, respectively. Section 2 shows that ICL for word analogy tasks can arise via modeling co-occurrence information using classical language models like continuous bag of words (CBOW) (violet represents relationship-specific nuisance tokens). Section 3 establishes the necessity of modeling positional information and blocked nuisance structure for ICL tasks, enabling pattern recognition and generalization to novel tokens (violet represents nuisance tokens). Section 4 presents scenarios where ICL fails, providing theoretical explanations that underscore the critical role of training data structure in enabling ICL in language models.
@@
Finally, we present two scenarios where ICL fails regardless of model architectures (see right of Figure 1). In the first scenario (left example), we consider a logic reasoning task that involves identifying and completing meta-patterns within sequences. Here, each training sequence repeats the pattern established by its starting tokens; the ICL task sequence then requires the model to recognize this meta-pattern of repetition and generalize it to a novel, unseen starting pattern. In the second scenario (right example), we examine a word analogy completion task in which relevant word pairs appear in the unstructured training sentences but are restricted to specific fixed positions. These findings, along with their empirical and theoretical explanations, underscore that LLMs require specific structures in the pre-training data to exhibit ICL ability.
@@
Summary of contributions. We (1) theoretically and empirically show that ICL for word analogy tasks with semantically related word pairs can arise from modeling co-occurrence patterns using CBOW, (2) prove that, to recognize token patterns and generalize them to novel tokens, ICL requires modeling positional information and blocked nuisance structure, and (3) present scenarios where ICL fails, highlighting the crucial role of training data structure for ICL.
Related work. Below, we highlight some of these studies and explain how our research aligns with, yet differs from, these approaches. We include a detailed discussion of related work in Appendix A. Numerous studies have connected ICL to classical methods, including gradient descent [5, 55, 17, 66, 3], Bayesian inference [56, 67, 14], and Newton’s method [20]. In contrast, our work links ICL to the continuous bag of words (CBOW) model, showing that ICL for word analogy tasks can be achieved by learning co-occurrence patterns. Several studies have examined the pre-training aspects of ICL, such as data distribution [37, 11, 30] and task diversity [42, 63]. By comparison, our work emphasizes the importance of co-occurrence, positional information, and training data structure for ICL to arise. Other research has explored ICL in specific data-generating processes, such as discrete functions [8] and autoregressive processes [45]. In contrast, our work centers on data characterized by semantically related word pairs and repeating token patterns.
@@
In this section, we focus on in-context learning (ICL) for word analogy tasks involving word pairs that frequently co-occur in training sentences; see Figure 1 (left). To motivate the discussion, we present two experiments using the LLaMA 2 model [51] involving countries (or US states) and their capital cities (see Appendix J for data sources). The prompts follow the format $c_{1}d_{1},c_{2}d_{2},\cdots,c_{6}d_{6},c_{7}$, where $c_{i}$ is a country (or US state) and $d_{i}$ is its capital city. In this scenario, we consider ICL successful if the model outputs $d_{7}$—the capital city of $c_{7}$—as the most likely token.
Experiment 1. We consider all 160 countries with a population exceeding one million in 2022. Among these countries, 31 have capital cities that are not their most populous cities, denoted by type A. The remaining 129 countries fall under type B. Each ICL prompt includes three type A countries among $c_{1},\cdots,c_{6}$ to emphasize that the desired relationship is (country)-(capital) rather than (country)-(largest city). Subsequently, we randomly generate 1,000 prompts, with 500 having a $c_{7}$ being a type A country and 500 having a $c_{7}$ being a type B country. The model’s ICL accuracy is $0.58$ for type A and $0.96$ for type B.
Experiment 2. We consider all 50 US states, among which 33 are of type A and 17 are of type B, similarly defined. Following the setup in Experiment 1, we generate prompts for these states. The ICL accuracy is $0.69$ for type A and $0.84$ for type B.
In both experiments, LLaMA 2 performs better on type B prompts (i.e., the capital city as the largest city). Since larger cities tend to appear more frequently as compared to smaller ones in the model’s pre-training data, this naturally raises the question: Can/does ICL with frequently co-occuring word pairs arise purely from modeling co-occurrence patterns?
ICL via classical non-transformer-based language models. We prove that, for word analogy tasks with frequently co-occurring word pairs, ICL can be achieved by modeling token co-occurrence—without needing positional encoding or attention mechanisms—using classical, non-transformer language models such as the continuous bag of words (CBOW) model [36]. (It does not imply that ICL in transformer-based models arises through learning co-occurrence patterns.) We utilize CBOW variant where each center word is modeled conditional on all other words in a sentence, not just neighboring words. Specifically, each word $w$ has center and context embeddings $u_{w}$ and $v_{w}$ of the same dimension. Given a sentence $x_{1}x_{2}\cdots x_{I}$, the $i$-th word ($x_{i}$) is distributed as $p(x_{i}=k\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{-i})\propto\exp((u_{k}^{\top}\sum_{j\neq i}v_{x_{j}})/(I-1))$, with $u_{w}$’s and $v_{w}$’s learned by minimizing cross-entropy losses across all positions.
@@
Roadmap of Section 2. In Section 2.1, we begin by considering a simple ICL task of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, where $(c_{i},d_{i})$ represents a frequently co-occuring word pair (e.g., a country and its capital city) and $i_{1},i_{2},\cdots,i_{\ell+1}$ are all distinct. The focus is to investigate whether a trained CBOW model can correctly output $d_{i_{\ell}}$. We also explore two other scenarios: ICL tasks of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$ in Section 2.2 (two connected word relationships), as well as $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell-1}}d_{i_{\ell-1}}c_{i_{\ell}}$ and $e_{i_{1}}f_{i_{1}}\cdots e_{i_{\ell}}f_{i_{\ell}}e_{i_{\ell+1}}$ (two disjoint word relationships) in Section 2.3. Section 2.4 concludes with synthetic experiments supporting the theory.
2.1 In-context learning on single-relationship word analogy tasks
@@
We investigate ICL in single-relationship word analogy tasks, where the training data contains only one type of relationship between frequently co-occurring word pairs. This task takes the form of $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, where $(c_{i},d_{i})$ pair represents a frequent co-occurrence, such as a country and its capital city. The vocabulary consists of $c_{1:K},d_{1:K},r_{1:L}$, where $r_{i}^{\prime}s$ represent other words (e.g., stop words). We first introduce Theorem 1, which states that ICL can arise if each sentence consists of exactly one $(c_{i},d_{i})$ pair, as long as the number of in-context examples ($\ell$) is not too large. To simplify calculations, we replace the cross-entropy loss with squared loss by removing the softmax activation and comparing outputs against the one-hot encoding of target words. The proof is in Appendix B.
Theorem 1 (ICL on single-relationship word analogy tasks).
@@
Let $K,L\geq S\geq 3$. Suppose each training sentence of length $S$ is generated by selecting one $(c_{i},d_{i})$ pair and $S-2$ distinct $r_{i}$’s uniformly at random. We train a CBOW model with the squared loss and a sufficiently large embedding dimension on these sentences. Given a prompt $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ with distinct $i_{k}$’s, the model correctly predicts $d_{i_{\ell+1}}$ if and only if
@@
As an example, when each training sentence contains exactly one country-capital pair (i.e., $(c_{i},d_{i})$), Theorem 1 says that a trained CBOW model will correctly predict $d_{i_{\ell+1}}$ (i.e., the capital city of $c_{i_{\ell+1}}$) given an ICL prompt of the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$, provided that the prompt length is not too large. Intuitively, this behavior is due to the presence of $c_{i_{\ell+1}}$ in the ICL prompt, leading the model to correctly predict $d_{i_{\ell+1}}$ given the frequent occurrences of the pair $(c_{i_{\ell+1}},d_{i_{\ell+1}})$ in the training data. However, when the prompt length is too large, the model will instead predict one of the $r_{i}$’s (see Theorem 1’s proof in Appendix B for more details). Moreover, if we let $L\rightarrow\infty$ and fix $K$ and $S$, the condition in Theorem 1 becomes $2\ell+1<K(S-1)^{2}/(S-2)^{2}$. This inequality trivially holds if the prompt length is set to be $S-1$ to match the length of the training sentences.
Furthermore, it is possible to adapt the proof of Theorem 1 to handle the case when each sentence comprises exactly two (not one) different $(c_{i},d_{i})$ pairs. In this case, letting $L\rightarrow\infty$ and fixing $K$ and $S$, the model correctly predicts $d_{i_{\ell+1}}$ given the same ICL prompt if and only if $2\ell+1<\frac{K(K-2)(S-1)^{2}}{(K-2)(S-2)(S-4)-K}$. This upper bound is strictly larger than $K(S-1)^{2}/(S-2)^{2}$: when each sentence contains exactly two $(c_{i},d_{i})$ pairs, ICL under the squared loss holds for longer prompts.
Experiments. To empirically verify Theorem 1 and its generalizations, we conduct experiments using the cross-entropy loss with $S=8$, $K=10$, $L=20$, and $\ell=3$. We explore multiple $(p_{0},p_{1},p_{2})$ values, where $p_{k}$ denotes the probability of having exactly $k$ pairs of $(c_{i},d_{i})$ in the sentence. For each $(p_{0},p_{1},p_{2})$ triple, we also introduce a more realistic setting where $c_{i}$ and $d_{i}$ do not always appear together by considering its corrupted version. In this setup, each $(c_{i},d_{i})$ pair has a 25% chance of being replaced with $(c_{i},r_{j})$ and a 25% chance of being replaced with $(d_{i},r_{j})$ for some $j\in[L]$. More details are provided in Appendix J.
Table 1: ICL on different single-relationship word analogy tasks, averaged over 10 repetitions, demonstrates stable, good performance across embedding dimensions ($d_{E}$), as Theorem 1 suggests. The corrupted setting also demonstrates excellent ICL ability under certain scenarios.
@@
Results. Table 1 displays the average accuracy for each scenario, calculated over 10 repetitions. Notably, when $(p_{0},p_{1},p_{2})$ is $(0,1,0)$ or $(0,0,1)$, ICL under the cross-entropy loss achieves zero accuracy, in contrast to perfect accuracy with the squared loss as shown in Theorem 1. We believe this difference in accuracy is an artifact of the loss functions used, although its relevance is limited by the fact that it is unlikely for every sentence to contain at least one $(c_{i},d_{i})$ pair, in reality. On the other hand, perfect ICL performance is observed in other settings (e.g., when the training sentences contain either zero, one, or two $(c_{i},d_{i})$ pairs) in both the clean and corrupted scenarios. For an in-depth comparison of ICL performance using both the squared and cross-entropy loss across various numbers of demonstration examples, see Appendix C.
2.2 In-context learning on dual-connected-relationship word analogy tasks
Building on the scenario that contains only a single type of relationship between frequently co-occurring word pairs, namely $(c_{i},d_{i})$, we now explore ICL on dual-connected-relationship word analogy tasks. Here, some words frequently co-occur with two different types of words in the training data, represented by the relationships $(c_{i},d_{i})$ and $(c_{i},e_{i})$. For instance, $c_{i}$ might represent a country, $d_{i}$ its capital city, and $e_{i}$ its currency. The vocabulary is comprised of $c_{1:K},d_{1:K},e_{1:K},r_{1:L}$, where $r_{i}$’s represent other words. Moreover, the corresponding ICL tasks take the form $c_{i_{1}}d_{i_{1}}\cdots c_{i_{\ell}}d_{i_{\ell}}c_{i_{\ell+1}}$ and $c_{i_{1}}e_{i_{1}}\cdots c_{i_{\ell}}e_{i_{\ell}}c_{i_{\ell+1}}$, where the model is expected to output $d_{i_{\ell+1}}$ and $e_{i_{\ell+1}}$, respectively. These can be regarded as task selection since the model should use the in-context examples to infer the tasks. We present Theorem 2, stating that a trained CBOW model can perform task selection if each sentence contains exactly two distinct $(c_{i},d_{i})$ or two distinct $(c_{i},e_{i})$ pairs with uniform probability. [^2] We can also theoretically show that ICL works (up to a certain number of training examples) in this scenario, but the calculations are extremely tedious. Therefore, we only present empirical evidence in Table 2.